OpenStudy (anonymous):

suqre root of 64 over squre root of 3 ? my only choices are 8 square root of 3 over 3 8 over 3 8 over square root of 3 or square root of 3 over 3 i don't understand how to work it out ?

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \frac{\sqrt{64}}{\sqrt{3}}\] \[\Large \frac{\sqrt{64}*\sqrt{3}}{\sqrt{3}*\sqrt{3}}\] \[\Large \frac{\sqrt{64*3}}{\sqrt{3*3}}\]

OpenStudy (anonymous):

\[\sqrt{64 * 3}\] is 24 right ?

jimthompson5910 (jim_thompson5910):

Oh I should have done it this way \[\Large \frac{\sqrt{64}}{\sqrt{3}}\] \[\Large \frac{8}{\sqrt{3}}\] \[\Large \frac{8*\sqrt{3}}{\sqrt{3}*\sqrt{3}}\] \[\Large \frac{8*\sqrt{3}}{\sqrt{3*3}}\] Hopefully that's a bit clearer The first way isn't wrong, it will just take more steps (than needed)

OpenStudy (anonymous):

 Okay, thank you. is that the right answer ^

jimthompson5910 (jim_thompson5910):

no, \[\Large \sqrt{3*3} =\sqrt{9} = 3\] this is all for the denominator

OpenStudy (anonymous):

Oh, okay. So  are possible answers ?